# Probabilistic Sampling of Balanced K-Means using Adiabatic Quantum Computing

### v1

@@ Abstract
Adiabatic quantum computing (AQC) is a promising quantum computing approach for discrete and often NP-hard optimization problems. Current AQCs allow to implement problems of research interest, which has sparked the development of quantum representations for many computer vision tasks. Despite requiring multiple measurements from the noisy AQC, current approaches only utilize the best measurement, discarding information contained in the remaining ones. In this work, we explore the potential of using this information for probabilistic balanced k-means clustering. Instead of discarding non-optimal solutions, we propose to use them to compute calibrated posterior probabilities with little additional compute cost. This allows us to identify ambiguous solutions and data points, which we demonstrate on a D-Wave AQC on synthetic and real data.
@@
1 Introduction
Clustering is a fundamental problem in machine learning and computer vision, extensively employed for the analysis and organization of large volumes of unlabeled visual data. It involves grouping similar objects based on their features, facilitating efficient data processing and unsupervised learning algorithms (Xu & Wunsch, 2005). The information extracted from clusters plays a vital role in various computer vision applications, including image classification (Lazebnik et al., 2006; Caron et al., 2018), segmentation (Coleman & Andrews, 1979; Dhanachandra et al., 2015), tracking (Keuper et al., 2020), and network training (Coates & Ng, 2012; Yang et al., 2016). However, the often NP-hard computational complexity of solving clustering problems often hinders their application to large-scale problems that require fast processing times (Xu & Wunsch, 2005).
In addition to this, the ambiguity of the task itself contributes to the absence of a unique definition for clustering (Xu & Wunsch, 2005), which arises from the range of possible optimization objectives. An example of this are different algorithms that can optimize for the most compact solution, an underlying data-generating process or just use local distances. One way to address this challenge is the use of confidence-based estimators and the sampling of multiple likely solutions. This approach helps in identifying well-assigned samples and uncovering ambiguities within the data (Xu & Wunsch, 2005).
Quantum algorithms that promise a considerable speedup over their classical counterparts and are inherently probabilistic have the potential to enable a new family of machine learning algorithms. By now, quantum machine learning algorithms approach tasks such as optimization of quadratic problems (Kadowaki & Nishimori, 1998), training of restricted Boltzmann machines (Dixit et al., 2021), and learning with quantum neural networks (Abbas et al., 2021). While the current quantum computers can only solve small-scale problems, they provide the basis to develop and test algorithms that can considerably increase the size of feasible problems in the future.
In our work, we propose to exploit the probabilistic nature of a quantum computer to sample multiple high-probability solutions of the balanced k-means problem at little additional cost. By formulating the k-means objective as a quadratic energy function, we embed the clustering task into the quantum-physical system of an AQC. By repeatedly measuring the quantum system, we sample high-probability solutions according to the Boltzmann distribution. Unlike previous approaches that only use the best solution and discard all other measurements, we utilize all samples to generate probabilistic solutions for the k-means problem, as shown in Figure 1. We recalibrate the samples of the clustering problem to address temperature mismatch (Pochart et al., 2021), estimating the posterior probability for each solution, which allows us to identify ambiguous points and provides alternative solutions. We demonstrate the algorithm on a D-Wave quantum computer and perform extensive experiments in simulation. However, our primary objective is to explore the potential and efficient utilization of quantum computing for machine learning. The primary contributions of our work are as follows:
@@ 1 Introduction
  • We propose a quantum computing formulation of balanced k-means clustering that predicts well-calibrated confidence values and provides a set of alternative clustering solutions.
@@
  • A reparametrization approach is used to compute posterior probabilities from samples that avoids exact tuning of the AQC sampling temperature.
@@
  • Extensive experiments on synthetic and real data show the calibration of our approach using simulation as well as the D-Wave Advantage 2 AQC prototype.
@@ 2 Related Work
With the availability of quantum computers to the general research community (Bunyk et al., 2014; McGeoch & Farré, 2021; McGeoch et al., 2022; Debnath et al., 2016), the research interest in finding applications for such systems has considerably increased. In this context AQC (Kadowaki & Nishimori, 1998; Bunyk et al., 2014) provides a well-tangible starting point, even though many applications need a complete reformulation considering the architectural differences of a quantum computer. Current applications for AQC include solving optimization problems for improved traffic flow (Neukart et al., 2017), robotic routing problems (Ohzeki et al., 2019), the design of molecules (Mulligan et al., 2020) or portfolio optimization in the finance sector (Mugel et al., 2021). Recently, the computer vision community has developed a strong interest in finding applications for quantum computing, which are related to hard permutation problems (Birdal et al., 2021) like tracking (Zaech et al., 2022) and graph-matching (Benkner et al., 2021).
@@
Clustering is a well-studied problem for quantum- as well as quantum-inspired algorithms (Aïmeur et al., 2007). Quantum clustering (Horn & Gottlieb, 2001) uses the Schrödinger equation to model the clustering problem, where cluster centers are defined as the minima of the corresponding potential function. Casaña-Eslava et al. (2020) extend this formulation with a probabilistic estimate of cluster memberships. These approaches use classical computation and gate-based quantum computers that are currently still on a small scale.
@@
Closest to this work, Arthur & Date (2021) present a balanced k-means clustering algorithm suitable for an AQC. While our underlying clustering algorithm follows the same approach, Arthur & Date (2021) discard all but the best measurement. Our approach, however, utilizes all information by employing AQC as a sampler to generate probabilistic solutions of the clustering problem, rather than only using the best solution in an optimization framework.
3 Theory
3.1 Quantum Computing
@@
Quantum computing is a fundamentally new approach, that utilizes the state of a quantum system to perform computations. In contrast to a classical computer, the state is probabilistic and described by its wave function, which enables the use fundamental properties of quantum systems like superposition and entanglement. Quantum algorithms that can efficiently run on these systems allow to solve previously infeasible problems, including the well-known prime factorization by Shor’s algorithm (Shor, 1999). In the following, the most important basics of quantum computing are explained.
@@
The qubit forms the basic building block of a quantum computer. In contrast to the classical bit, which is always in either of its two basis states |$\ket{0},\ket{1}$, the qubit can exist in a state of superposition, which is a linear combination of any two basis states
|$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}.$ (1)
The two complex-valued factors $\alpha,\beta$ describing the superposition are called amplitudes.
@@
To perform meaningful operations, a quantum computer operates on a system of multiple qubits. If the qubits are independent of each other, the joint state of the system can be computed as the tensor-product of all involved qubit states. If the qubits of a system are entangled (Einstein et al., 1935; Schrödinger, 1935), it is is not possible to decompose the joint state into the tensor-product of each separate qubit state. Therefore, measuring the state of one qubit in a entangled system, affects the state of the other qubits. This is a fundamental property of quantum mechanics and plays a crucial role for the capabilities of quantum computing.
@@
During computation on the quantum computer, the state of the system can be any valid superposition of basis states. Nevertheless a measurement of the system always results in a single basis state, which is referred to as wave-function collapse (Wilde, 2017). The probability of measuring a state is the respective squared amplitude. In the single qubit case as described in equation 1, this corresponds to
@@
3.2 Adiabatic Quantum Computing.
AQC, which is used in this work, is a quantum computing paradigm where the state of a quantum system is modified by performing an adiabatic transition. Current hardware implementations such as the D-wave systems (Bunyk et al., 2014) follow this approach and implement a Quantum Annealer (QA) to solve quadratic unconstrained binary optimization (QUBO). They are based on the Ising model (Ising, 1925; Kadowaki & Nishimori, 1998), which describes the configuration of a set of interacting particles that all carry an atomic spin $\sigma_{i}$.
@@
The lowest energy state is found by following the adiabatic theorem (Born & Fock, 1928). Starting with an initial system in its ground state described by an Hamiltonian $H_{0}$, the adiabatic theorem states that during a sufficiently slow change of the Hamiltonian the system never leaves its ground state. The change of the Hamiltonian
@@
is called an adiabatic transition. The transition is performed over the annealing-time $T_{a}$ and allows to solve an optimization problem with the Hamiltonian $H_{T}$.
@@
$p(\mathbf{\sigma})=\frac{\mbox{exp}[\nicefrac{{-H(\mathbf{\sigma})}}{{T}}]}{\sum_{\mathbf{\sigma}^{\prime}}\mbox{exp}[\nicefrac{{-H(\mathbf{\sigma}^{\prime})}}{{T}}]},$ (5)
@@
where $p(\mathbf{\sigma})$ describes the probability of finding the system in state $\mathbf{\sigma}$ and $\mathbf{\sigma}^{\prime}$ are all possible states. In this work, we use this property, as it allows us to sample from the Boltzmann distribution corresponding to the energy-based model of clustering.
3.3 Energy-Based Models
Energy-based models are probabilistic models that use an energy function $E(\mathbf{x})$ to describe the probability of each system state. By assigning an energy value to each system state, the probabilities are Boltzmann distributed
$p(\mathbf{x})=\frac{1}{A}\mbox{exp}[\nicefrac{{E(\mathbf{x})}}{{T}}],$ (6)
with the temperature $T$ as a free parameter in this model and the normalization constant $A$ ensuring that the values are valid probabilities. Our model follows this approach and represents the clustering problem as an energy-based formulation.
One challenging task to use energy-based models remains in sampling the system according to the Boltzmann distribution as this is NP-hard for the quadratic Ising-model formulation (Pochart et al., 2021). While this is usually approximated on classical hardware with the corresponding computational overhead, AQC provides a direct way to sample from a physical system that follows the Boltzmann distribution (Pochart et al., 2021).
3.4 Clustering
Clustering is one core task required in many unsupervised machine learning algorithms. It has the objective to group a set of points $\mathbf{X}$ into disjoint clusters $\{c_{1},...,c_{K}\}$, where each cluster contains points that are similar in the feature space. A design parameter is the definition of the similarity metric, which leads to a wide range of clustering algorithms available. The popular and simple k-means algorithm (Forgy, 1965; Lloyd, 1982) optimizes the quadratic distance to a centroid. Such distance-based metrics assume clusters to be compact, while other approaches such as dbscan (Ester et al., 1996) define features as similar if they come from a contiguous region of high density.
In this work, balanced clustering is investigated, where a defined target cluster size $s_{k}$ is required. This approach forms the basis of many AQC-based algorithms (Benkner et al., 2020; Birdal et al., 2021; Zaech et al., 2022) and has a wide range of further applications such as secret key generation in cryptography (Hong et al., 2017), energy-efficient data aggregation in wireless sensor networks (Liao et al., 2013), or data cleaning (Fisher et al., 2015).
Uncertainty estimation in clustering and machine learning aims at quantifying the confidence that predictions made by a model correspond to the underlying ground-truth. The confidence scores can be used to eliminate uncertain samples from the solution, to select a set of possible hypotheses or to find the right parameters of the algorithm itself (Casaña-Eslava et al., 2019).
Uncertainty estimates can be generated separately for each data-point by evaluating the likelihood function $p(\mathbf{X}|{\hat{Z}})$, or by computing the posterior probability $p({\hat{Z}}|\mathbf{X})$ that describes the probability of the whole clustering solution. As the former can be evaluated if the data generating model is defined, it can be combined with most clustering approaches. Nevertheless, as the likelihood function is not a probability distribution of the cluster assignments, it does not provide a calibrated prediction. In contrast to this, the posterior probability $p({\mathbf{Z}}|\mathbf{X})$ directly represents the probability that the estimated assignment $\mathbf{Z}$ corresponds to the ground-truth. While this provides an interpretable result, it is often infeasible to compute as all possible assignments need to be evaluated in order to find the normalizing constant.
The set of possible clustering solutions together with their posterior probabilities provides valuable information that can be utilized for a range of low- as well as high-level reasoning tasks. On a low level, the probability of the best solutions can be used to identify the right number of clusters. If a number different than the number of the data-generating process is chosen, additional ambiguity is introduced. For a larger number during clustering than present in the data, the strong overlap between the corresponding distributions induces high ambiguity, while an insufficient number of clusters spreads points inside a single cluster far apart, which reduces the associated probability.
For high-level applications, calibrated samples of possible clustering solutions can provide additional information in multi-object tracking (Chiu et al., 2021). Following the AQC framework, tracking can be implemented as a clustering problem with additional constraints that represent the temporal relation between points (Zaech et al., 2022). Each cluster then corresponds to one object in the video, with each point representing it at a different timestep. The feature used to define clusters can either be visual similarity from re-identification(Hirzer et al., 2011; Paisitkriangkrai et al., 2015; Zheng et al., 2019) features or spatial similarity(Bewley et al., 2016). Different solutions thus represent different tracks through time, where ambiguities can be generated by occlusions or crossing paths. In a larger system such as an autonomous vehicle, knowing all likely candidate solutions can allow for predicting multimodal candidates for future trajectories (Jiang et al., 2023; Deo & Trivedi, 2018) and can help to evaluate the risk of taking any action. In a similar approach, feature matching can be formulated (Birdal et al., 2021) and ambiguous candidates can be discarded during 3d reconstruction or camera pose estimation.
3.5 Clustering as QUBO

### v2

# heightadjust=object

@@ Abstract
Adiabatic quantum computing (AQC) is a promising approach for discrete and often NP-hard optimization problems. Current AQCs allow to implement problems of research interest, which has sparked the development of quantum representations for many computer vision tasks. Despite requiring multiple measurements from the noisy AQC, current approaches only utilize the best measurement, discarding information contained in the remaining ones. In this work, we explore the potential of using this information for probabilistic balanced k-means clustering. Instead of discarding non-optimal solutions, we propose to use them to compute calibrated posterior probabilities with little additional compute cost. This allows us to identify ambiguous solutions and data points, which we demonstrate on a D-Wave AQC on synthetic tasks and real visual data.
1 Introduction
@@
Clustering and uncertainty quantification (UQ) are fundamental problems in machine learning and computer vision. Clustering, the task of grouping objects based on the similarity of their features, plays a pivotal role in analysis and organization of vast quantities of unlabeled visual data with applications including image classification [51, 16, 65], segmentation [23, 32], tracking [48], and network training [75, 22, 76]. UQ, on the other hand, aims to assess the trustworthiness of predictions and accounts for the impact of data variability. As a crucial element for certifying confidence in results, UQ is indispensable in clustering, especially given the often ambiguous nature of data partitioning [41].
Clustering typically involves solving for an optimal assignment of data points to (latent) centroids, while a full Bayesian UQ approach strives to accurately depict the posterior distribution. Unfortunately, both of these problems are known to be notoriously difficult, classified within the NP complexity class [^1] Computing marginal probabilities on a Bayesian network is, basically, a counting problem that can become arbitrarily hard to solve. [24, 26, 55]. These challenges within the Turingian modes of computation have motivated: (i) continuous relaxations in clustering [78] and (ii) the advent of approximate Bayesian inference methods for UQ, including the EM algorithm [30], variational inference [13], and sequential MCMC [29].
In this paper, departing from the above-mentioned approximations or continuous relaxations, we take a completely different approach and harness the upcoming and novel computational paradigm of quantum computers to tackle the challenging task of clustering with calibrated uncertainty quantification. Specifically, we focus on the widely used balanced K-means [52, 54, 56], which is an iterative algorithm that first assigns each data point to a cluster followed by an update of cluster centroids.
Despite their early state, quantum machine learning algorithms have approached tasks such as optimization of quadratic problems [47], training of restricted Boltzmann machines [33], and learning with quantum neural networks [1]. While the current quantum computers can only solve small-scale problems, they provide the basis to develop and test algorithms that can considerably increase the size of feasible problems in the future. In our work, we interpret K-means clustering in an energy-based framework and aim to model the posterior distribution by sampling multiple high-probability binary assignments from it on actual quantum hardware at little additional cost. We specifically use the quantum annealer (QA) of D-Wave [58], which follows the concept of adiabatic quantum computing (AQC) [10] and avoids simulating costly Markov chains as in classical computers. To achieve this, we first formulate the k-means objective as a quadratic energy function over binary variables and embed the clustering task into the quantum-physical system of D-Wave. By repeatedly measuring the quantum system, we sample high-probability solutions according to the Boltzmann distribution. Unlike previous approaches that only use the best solution and discard all other measurements, we utilize all samples to generate probabilistic solutions for the k-means problem, as shown in Figure 1. Due to temperature mismatch [68], AQCs sample from a modified posterior instead of the true posterior, which needs to be re-calibrated for the task at hand. We do so by estimating the posterior probability for each solution. This allows us to identify ambiguous points and provide alternative solutions. By leveraging QA, our approach respects the binary nature of the assignment problem.
We demonstrate our algorithm on a D-Wave quantum computer and also perform extensive experiments in simulation. On real data, we show that our approach can be utilized for distributions that do not strictly follow the initial assumptions and that it is suitable for identifying ambiguous images. Our primary contributions are:
@@
  • A quantum computing formulation of balanced k-means clustering that predicts calibrated confidence values and provides a set of alternative clustering solutions.
@@
  • A reparametrization approach used to calibrate posterior probabilities from samples that avoids exact tuning of the AQC sampling temperature.
@@
  • Extensive experiments on synthetic and real data showing the calibration of our approach both on simulations and on the D-Wave Advantage 2 QA prototype.
Developing better heuristics for the NP-hard challenges of clustering and UQ presents significant difficulties for classical approaches. However, we are optimistic that the ongoing progress in quantum computing technology will progressively enhance the effectiveness of our algorithms.
@@ 2 Related Work
With the availability of quantum computers to the general research community [15, 57, 58, 28], the research interest in finding applications for such systems has considerably increased. In this context AQC [47, 15] provides a well-tangible starting point, even though many applications need a complete reformulation considering the architectural differences of a quantum computer. Current applications for AQC are optimization tasks in different fields [64, 66, 62, 61], including quantum computer vision [12], where a strong interest in finding quantum computing formulations has developed. These are often related to hard permutation problems [7, 12, 6] like tracking [77], graph, shape and point matching [6, 9, 60] or training binary (graph) neural networks [49]. In this context, Birdal et al. [12] evaluate the k-best solutions, however, without the energy-based formulation, only little improvement is achieved. Another tasks of interest for the community is model fitting to find camera parameters [34] or separating motion components [3], which can be formulated as a quantum-computing problem [74, 20, 3, 34, 19, 37] instead of consensus maximization.
Clustering.
@@
Clustering is a well-studied problem for quantum- as well as quantum-inspired algorithms [2]. It groups a dataset ${X}$ into disjoint clusters ${c_{1},...,c_{K}}$, with each cluster containing points similar in the feature space. For example, k-means algorithm [39, 53] utilizes quadratic distances, favoring compactness, whereas dbscan [36] identifies local high-density regions. Quantum clustering methods use either the paradigm of AQC [5, 4, 65, 50] or circuit-based quantum computing such as Quantum clustering [43] that models clustering through the Schrödinger equation, where cluster centers are defined as the minima of its potential function. Casaña-Eslava et al. [18] extend this formulation with a probabilistic estimate of cluster memberships. These approaches use classical computation and circuit-based quantum computers that are fundamentally different from AQC and currently still on a smaller scale.
@@
Closest to this work, Arthur and Date [4] present a balanced k-means clustering algorithm with predefined target size $s_{k}$ suitable for an AQC and Nguyen et al. [65] use a similar formulation to cluster visual features. While our underlying algorithm follows a similar approach as [4, 65], they discard all but the best measurement. Our approach, however, utilizes all information by employing AQC as a sampler to generate probabilistic solutions of the clustering problem, rather than only using the best solution in an optimization framework. We particularly consider balanced clustering, as it forms the basis of several AQC algorithms [7, 12, 77] in computer vision.
Uncertainty quantification in clustering.
Knowing the set of high-quality solutions and their confidence offers valuable insights for both low- and high-level tasks. E.g. at a low level, the probability of top solutions can help to determine the correct cluster count. Choosing a different number than the actual data process introduces ambiguity [17]: too many clusters cause overlap, while too few clusters spread points, reducing associated probability.
For high-level tasks, calibrated clustering solutions have the potential to enhance matching problems like multi-object tracking [21]. Following the AQC framework, tracking becomes a clustering problem with additional constraints that represent the temporal relation between points [77], where each cluster corresponds to one track. The feature defining clusters can be visual similarities from re-identification [42, 67, 79] or spatial similarity [8]. The set of solutions models different trajectories due to occlusions or intersecting paths. In complex systems like autonomous vehicles, understanding multiple probable solutions helps to predict multimodal future trajectories [46, 31], to assess risks when taking actions. In a similar approach, AQC is used for feature matching [12] and uncertainty estimation potentially allows to discard ambiguous candidates during 3d reconstruction.
3 Adiabatic Quantum Computing
@@
Quantum computing is a fundamentally new approach, that utilizes the state of a quantum system to perform computations. In contrast to a classical computer, the state is probabilistic and described by its wave function, which enables the use fundamental properties of quantum systems like superposition and entanglement. Quantum algorithms that can efficiently run on these systems allow to solve previously infeasible problems, including the well-known prime factorization by Shor’s algorithm [71]. In the following, the most important basics of quantum computing are explained.
@@
Qubits are the building block of a quantum computer. In contrast to the classical bit, which is in either of its two basis states |$\ket{0},\ket{1}$, the qubit can exist in a state of superposition, which is a linear combination of any two basis states |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$. The two complex-valued factors $\alpha,\beta$ describing the superposition are called amplitudes.
@@
To perform meaningful operations, a quantum computer operates on a system of multiple qubits. If the qubits are independent of each other, the joint state of the system can be computed as the tensor-product of all involved qubit states. If the qubits of a system are entangled [35, 70], it is is not possible to decompose the joint state into the tensor-product of each separate qubit state. Therefore, measuring the state of one qubit in a entangled system, affects the state of the other qubits. This is a fundamental property of quantum mechanics and plays a crucial role for the capabilities of quantum computing.
@@
During computation on the quantum computer, the state of the system can be any valid superposition of basis states. Nevertheless, a measurement of the system always results in a single state of the measurement basis, which is referred to as wave-function collapse [73]. The probability of measuring a state is the respective squared amplitude. In the single qubit case, this corresponds to
@@
Adiabatic Quantum Computing.
AQC, which is used in this work, is a quantum computing paradigm where the state of a quantum system is modified by performing an adiabatic transition. Current hardware implementations such as the D-wave systems [15] follow this approach by implementing QA to solve quadratic unconstrained binary optimization (QUBO). They are based on the Ising model [45, 47], which describes the configuration of a set of interacting particles that all carry an atomic spin $s_{i}$.
@@
The lowest energy state is found by following the adiabatic theorem [14]. Starting with an initial system in its ground state described by a Hamiltonian $H_{0}$, the adiabatic theorem states that during a sufficiently slow change of the Hamiltonian the system never leaves its ground state. The change of the Hamiltonian is called an adiabatic transition and is often performed with a linear schedule
@@
over the annealing-time $T_{a}$ and allows to solve an optimization problem with the Hamiltonian $H_{T}$.
@@
$p({s})={\mbox{exp}[\nicefrac{{-H({s})}}{{T}}]}/{\sum_{{s}^{\prime}}\mbox{exp}[\nicefrac{{-H({s}^{\prime})}}{{T}}]},$ (4)
@@
where $p({s})$ describes the probability of finding the system in state ${s}$ and ${s}^{\prime}$ are all possible states. In this work, we use this property, to sample from the Boltzmann distribution corresponding to the energy-based model of clustering.
4 Balanced Quantum K-Means
In the following, we define an energy-based model (EBM) that directly connects to AQC. Based on this, we derive the energy function for k-means and demonstrate how an AQC can sample from the corresponding distribution.
EBM Inference.
We use the variable $Z\in\mathbb{Z}^{N}_{2}$, to represent our binary parameters. The posterior density of interest follows from the Bolzmann distribution as:
$\pi(Z)\vcentcolon=p(Z|X)\propto\exp(-E(Z,X)),$ (5)
where $X$ denotes the observations, a.k.a. the collection of all specified data and $E$ is called the potential energy:
$E(Z,X)\vcentcolon=-(\log p(X|Z)+\log p(Z)).$ (6)
Following [11, 12], we consider the probabilistic inference, where we will be interested in the following quantities:
I Maximum a-posteriori (MAP):
$\hat{Z}=\operatorname*{arg\,max}\limits_{Z\in\mathbb{Z}^{N}_{2}}\log p(Z|X)$ (7)
where $\hat{Z}$ denotes the entirety of the sought parameters.
II The full posterior distribution: $p(Z|X)\propto p(X,Z)$.
Both of these problems are very challenging and cannot be directly addressed by standard methods such as gradient descent (problem I) or standard MCMC methods (problem II). The difficulty in these problems is mainly originated by the fact that the posterior density is non-log-concave (i.e. the negative log-posterior is non-convex) and any algorithm that aims at solving one of these problems should be able to operate in the particular space of binary variables. Nevertheless, The MAP estimate is often easier to obtain via a Quantum annealer and useful in practice [12, 7]. On the other hand, samples from the full posterior can provide important additional information, such as uncertainty. Not surprisingly, the latter is a much harder task, especially considering the discrete nature of our problem. Unfortunately, available QAs such as D-Wave [58] are not directly capable of tackling the second problem since they sample from a modified posterior and not the true one. Our work fills this gap for the task of balanced K-means clustering using quantum annealing as a direct way to sample from a physical system that follows the Boltzmann distribution [68] and recomputes the posterior from the found solutions to ensure callibration.
Clustering as QUBO.